OpenStudy (anonymous):

The length of a rectangle is twice its width. If the perimeter of the rectangle is 30yd , find its area

OpenStudy (anonymous):

area is 50

OpenStudy (anonymous):

should be exact 50 yd squared

hero (hero):

Juanita, you should probably show your work in addition to providing the answer. . Just posting answers is not very helpful.

OpenStudy (anonymous):

\[2(x) + 2(2x) = 30\] where x is the shorter side. \[6x = 30\] \[x = 5\] so \[x \times 2x = 5 \times 10 = 30\]

OpenStudy (anonymous):

I mean 50 sorry, the final answer is 50 Fail

hero (hero):

That's why I avoid using the equation editor when it isn't necessary.

OpenStudy (anonymous):

I'm new so I just wanted to give it a whirl XD.

hero (hero):

I only use it for clarification purposes.